OpenStudy (anonymous):

how many years older will you be 1.70 billion seconds from now ( assume a 365 days a year)

OpenStudy (aum):

1 year = 365 days = 365 * 24 hours = 365 * 24 * 60 minutes = 365 * 24 * 60 * 60 seconds. 1.7 billion seconds = 1.7 x 10^9 seconds = 1.7 x 10^9 / (365 * 24 * 60 * 60) years = ?

OpenStudy (anonymous):

From what I remember, 30 years old is 1 billion seconds, so at least you know it's older than 30 :P, but aum did the work for ya, so you just have to plug in the numbers in a calculator :).